Circular No. 8636

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2005 (2003 EL_{61}) 2

M. E. Brown, California Institute of Technology, on behalf of the adaptive-optics team at Keck Observatory, reports the discovery of a second satellite of the transneptunian object 2003 EL_{61} from K′-band images taken with the Laser Guide Star Adaptive Optics system at the Keck II telescope on Mar. 1, May 27, and June 29. The satellite is 4.6 ± 0.5 mag fainter than the primary. A preliminary circular orbit suggests a 34.1-day period with a semimajor axis of 39300 km, inclined by 40^{o} from the larger outer satellite, S/2005 (2003 EL_{61}) 1 (cf. IAUC 8577), though additional observations are required to confirm the orbit. Additional information can

be found at website http://www.gps.caltech.edu/~mbrown/2003EL61

NOVA IN THE LARGE MAGELLANIC CLOUD 2005 E. O. Waagen, AAVSO, reports that B. Allen’s position (cf. IAUC

8635) was only an estimation, adding that a precise position was measured by P. Nelson (Ellinbank, Victoria, Australia) from a CCD image taken on Nov. 27.583 UT: α = 5^{h}10^{m}32^{s}.68, δ = −69^{o}12′35′′.7 (equinox 2000.0).

SUPERNOVA 2005ky D. A. Howell, University of Toronto, on behalf of the Supernova Legacy

Survey (http://xxx.lanl.gov/abs/astro-ph/?0510447), reports the dis-

covery of a supernova at magnitude i′(AB) = 22.3 on Nov. 24.3 UT on Megacam images obtained with the Canada-France-Hawaii Telescope. SN 2005ky, which had brightened to i′(AB) = 21.0 by Dec. 1.3, is located at R.A. = 2^{h}24^{m}36^{s}.254, δ = −4^{o}10′54′′.94 (equinox 2000.0), which is 0′′.33 west and 3′′.40 south of the center of its host galaxy. A spectrogram of 2005ky, obtained by M. Sullivan and R. S. Ellis on Nov. 30.3 with the Keck I telescope (+ LRIS), indicates that this is a type-Ia supernova about one week before maximum light. A redshift of z = 0.148 was determined from host-galaxy Ca H and K absorption. The supernova shows a blue wing on the Si II 635.5-nm feature that is a possible indication of high-velocity features. The expansion velocities are relatively high for this epoch: 15000 km/s determined from Si II 635.5-nm and 27000 km/s determined from Ca II H and K and the Ca II infrared triplet. There is no apparent feature at 580 nm, and the blue part of the S II ‘W’ at 540 nm is enhanced relative to the red absorption. Strong Na D from the host galaxy indicates that there may be extinction along the line-of-sight to this supernova.

2005 December 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green